# Density-based User Representation using Gaussian Process Regression for Multi-interest Personalized Retrieval

### v4

@@ 4.4 Item Embedding Pre-training
Following Pal et al. (2020), we assume that item embeddings are fixed and precomputed: this ensures rapid computation and real-time updates at serving time. For item embedding pre-training, we use extreme multi-class classification, a method widely used in prior work on sequential recommendation (Cen et al., 2020; Kang & McAuley, 2018; Hidasi et al., 2016): given a training sample $(u_{i},v_{j})$, we first compute the likelihood of $u_{i}$ interacting with $v_{j}$, i.e., $p(v_{j}|u_{i})=\text{exp}(\mathbf{u}_{i}^{\intercal}\mathbf{v}_{j})/\sum_{v^{\prime}\in\mathcal{V}}\text{exp}(\mathbf{u}_{i}^{\intercal}\mathbf{v}^{\prime})$, where $\mathbf{u}_{i}$ and $\mathbf{v}_{j}$ are embeddings of $u_{i}$ and $v_{j}$. Our objective is to maximize the log-likelihood of a user interacting with their items. Moreover, we want to ensure that item embeddings align with their categories since categories explicitly indicate a user’s interests. We capture item-category information by computing the likelihood that an item belongs to a category in a similar way: $p(c_{k}|v_{j})=\text{exp}(\mathbf{v}_{j}^{\intercal}\mathbf{c}_{k})/\sum_{c^{\prime}\in\mathcal{C}}\text{exp}(\mathbf{v}_{j}^{\intercal}\mathbf{c}^{\prime})$.
The overall objective for pre-training combines the two negative log-likelihoods using a scaling factor $\gamma$:
@@
where $\mathcal{V}_{u_{i}}$ is the set of items that $u_{i}$ interacted with and $\mathcal{C}_{v_{j}}$ is the set of categories that $v_{j}$ belongs to. The full item embedding matrix $\mathbf{V}$ is jointly learned and fixed after the pre-training phase. Notice that we do not use the user and category embeddings after pre-training, and other schemes for computing item embeddings are possible. For example, an alternative approach is to use item co-occurrence information to derive item embeddings only, without user or category embeddings. The full architecture for GPR4DUR is depicted in Figure 2.
@@ 5 Offline Experiments
We next evaluate our GPR4DUR on three real-world datasets, and compare our DUR technique with other state-of-the-art methods for multi-interest retrieval.
@@
The higher the better normal-$\Uparrow$ The higher the better normal-$\Uparrow$ The lower the better normal-$\Downarrow$ The higher the better normal-$\Uparrow$
@@ 5.1 Datasets
We use three widely studied datasets: Amazon (He & McAuley, 2016), MovieLens (Harper & Konstan, 2015), and Taobao (Zhu et al., 2018). We adopt the same setting as previous work (Li et al., 2020; Wang et al., 2019), filtering out items with fewer than ten appearances in the dataset, and users with fewer than 25 item interactions. This ensures users have sufficient history to indicate multiple interests. Each interest corresponds to a category: for Amazon, we use the single principle category of each item; for MovieLens, the 18 movie genres are categories—each movie can belong to multiple categories; and for Taobao, we cluster items into 20 categories using Ward clustering on the pretrained item embeddings, a technique shown to be effective by Pal et al. (2020). We use the same categories during training and inference. Dataset statistics shown in Table 5 (See Appendix). An ablation analysis of various clustering algorithms and numbers of clusters is shown in Table 6 (See Appendix).
@@ 5.2 Experiment Setup
Most prior work on recommendation and retrieval evaluates model performance on generalization to new items per user. Instead, similar to recent work (Liang et al., 2018a; Cen et al., 2020), we assess model performance on a more challenging task: its ability to generalize to new users. We split users into disjoint subsets: training users ($\mathcal{U}^{\text{train}}$), validation users ($\mathcal{U}^{\text{val}}$), and test users ($\mathcal{U}^{\text{test}}$) at a ratio of 8:1:1. We treat the last 20% of each user’s full interaction sequence as a holdout set for evaluation, and the first 80% as a history set used to fit the GPR model. We cap the history length to 60, 160, and 100 for the three datasets, aligning with the average count of user interactions per dataset.
For the item embedding pre-training phase, we train the recommendation backbones using the history set of all training users and tune the parameters based on the performance on the holdout set for validation users. The maximum number of training iterations is 100,000. Early stopping is used if performance on the validation set does not improve for 50 successive iterations. We tune GPR hyperparameters by fitting the GP regressor to the history set of all training and validation users, to learn their user representations, then tune the parameters (i.e., kernel $k$, variance $\sigma$) using the holdout set. To compare all methods fairly, we report metrics on the holdout set for all test users.
@@ 5.3 Metrics
Metrics used in prior work on sequential recommendation are not well-suited to assess performance in our multi-interest retrieval task, for several reasons: (i) Conventional metrics, such as precision and recall, often employed in multi-interest research, do not adequately quantify whether an item list reflects the full range of a user’s (multiple) interests. A model may primarily recommend items from a narrow range of highly popular categories and still score high on these metrics while potentially overlooking less popular or niche interests. (ii) Metrics like precision and recall are overly stringent, only recognizing items in the recommendation list that appear in the holdout set. We argue that credit should also be given if a similar, though not identical, item is recommended (e.g., Iron Man 1 instead of Iron Man 2). This requires a soft version of these metrics. (iii) Multi-interest retrieval systems should expose users to niche content to address their diverse interests. However, conventional metrics may neglect the item perspective, potentially underserving users with specialized interests. Consequently, we propose the use of the following four metrics that encompassing the aforementioned factors.
@@
This metric is similar to subtopic-recall (Zhai et al., 2003), which directly measures whether the model can comprehensively retrieve all user interests reflected in the holdout set. The higher the value of this metric the better:
@@
$\text{IR}@k$ $=\frac{1}{|\mathcal{U}^{\text{test}}|}\sum_{u\in\mathcal{U}^{\text{test}}}\frac{\sum\limits_{c\in\mathcal{C}(\mathcal{V}^{\text{d}}_{u})}\max\limits_{v_{i}\in\mathcal{V}^{\text{d}}_{u},v_{j}\in\mathcal{R}^{1:k}_{u}}S(v_{i},v_{j})}{|\mathcal{C}(\mathcal{V}^{\text{d}}_{u})|},$ (8)
s.t. $\mathcal{C}(v_{i})=\mathcal{C}(v_{j})=c,$
where $S(v_{i},v_{j})$ is the cosine similarity between item $v_{i}$ and $v_{j}$. To obtain ground-truth similarities between items, and to mitigate the influence of the chosen pre-trained model, we pretrain the item embeddings using YoutubeDNN (Covington et al., 2016a) with a higher dimension size ($d=256$) to compute a uniform $\mathbf{S}_{i,j}$ for any backbone. A higher value of this metric is better.
@@
$\text{ED}@k$ $=\frac{1}{|\mathcal{U}^{\text{test}}|}\sum\limits_{u\in\mathcal{U}^{\text{test}}}||\epsilon^{*}_{u}-\epsilon_{u}^{1:k}||_{2}^{2},$
s.t., $\epsilon^{*}_{u,c}=\frac{\sum_{v\in\mathcal{V}^{\text{d}}_{u}}\mathbbm{1}_{{c\in\mathcal{C}(v)}}}{\sum_{v\in\mathcal{V}^{\text{d}}_{u}}|\mathcal{C}(v)|},\epsilon^{1:k}_{u,c}=\frac{\sum_{v\in\mathcal{R}^{1:k}_{u}}\mathbbm{1}_{{c\in\mathcal{C}(v)}}}{\sum_{v\in\mathcal{R}^{1:k}_{u}}|\mathcal{C}(v)|}.$ (9)
@@
5.4 Methods Studied
We study the following nine methods from four categories.
(i) Heuristic Methods $\clubsuit$ : Random recommends random items, MostPop recommends the most popular items.
[…]
(iii) MUR Methods normal-$\spadesuit$ : MIND (Li et al., 2019) designs a multi-interest extractor layer based on the capsule routing, which is applicable for clustering past behaviors and extracting diverse interests. DGCF (Wang et al., 2020a) investigates the fine-grained user intention to enhance the ability of collaborative filtering by disentangling multiple latent factors for user representation. ComiRec (Cen et al., 2020) is one of the SOTA methods that captures multiple interests from user behavior sequences with a controller for balancing diversity. CAMI (Liu et al., 2022) uses a category-aware multi-interest model to encode users as multiple preference embeddings to represent user-specific interests.
[…]
@@ 5.4 Methods Studied
Our overall performance comparison addresses two central research questions (RQs): whether our proposed method, GPR4DUR, offers superior retrieval performance for users with multiple interests (RQ1); and whether it induces appropriate item-sided exposure w.r.t. both popular and niche interests (RQ2).
To assess RQ1, we evaluate the effectiveness of the retrieval phase. As shown in Table 1, GPR4DUR outperforms almost all baselines across all datasets w.r.t. Interest Coverage (IC@$k$) and Interest Relevance (IR@$k$), demonstrating high degrees of retrieval coverage and relevance, respectively. Specifically, on the Amazon dataset, GPR4DUR achieves the highest performance in 4 out of 6 interest metrics, in MovieLens it leads in 5 out of 6 metrics, and in Taobao, it excels across all 6 interest metrics. This shows that GPR4DUR covers a wide range of user interests while simultaneously maintaining excellent relevance.
For RQ2, our objective is to ascertain whether item exposure is suitably balanced. We measure this using the Exposure Deviation (ED@$k$) and Tail Exposure Improvement (TEI@$k$) metrics. Lower values of ED@$k$ suggest more satisfying category exposure, while higher values of TEI@$k$ are indicative of enhanced exposure in the long tail of item categories. GPR4DUR is highly effective on these metrics, consistently performing well across all datasets in most cases, and validating its ability to provide an optimal level of category exposure. Notably, GPR4DUR achieves the best results in eight of nine TEI@$k$ metrics, indicating its superior ability to generate exposure to niche categories/interests. We point out, however, that all TEI@$k$ values are negative, which suggests that none of the methods we tested improve exposure for niche categories relative to the exposure in the user holdout sets. This is explained by the well-known popularity bias inherent in most recommendation methods, and suggests that further work is required on diversification strategies to mitigate such effects.
While our primary emphasis is on the retrieval phase, we also present model comparisons for the ranking task. As illustrated in Table 2, GPR4DUR demonstrates competitive performance on the traditional relevance metrics relative to the baseline models. This further validates the strength of our proposed method.
@@
(RQ4) We underscored the importance of efficiency for a good user representation in Sec. 1. To shed light on this, we examine the IC@20 and IR@20 performance of various methods on the Amazon dataset. As shown in Fig. 4, GPR4DUR is very effective w.r.t. interest coverage even when operating with low-dimensional embeddings (i.e., $d=8$ and $d=4$). By contrast, the performance of the SUR and MUR methods degrades as the dimensionality decreases. GPR4DUR consistently outperforms other methods w.r.t. interest relevance across all dimensionalities examined. We note that lower dimensions facilitate higher interest relevance due to increased cosine similarity (Eq. 8), explaining the inverse correlation between IR@20 and dimension size in Figure 4 (right).
@@ 6 Online Simulation
Data Preparation. We assume $|\mathcal{C}|=10$ interest clusters, each represented by a $d$-dim ($d=32$) multivariate Gaussian. We randomly select a (user-specific) subset of these interests as the ground-truth interest set for each user. We set $|\mathcal{U}|=1000$ and $|\mathcal{V}|=3000$, with $300$ items in each interest cluster. Each item belongs to a single cluster and its embedding is sampled from the corresponding interest distribution. Each user is modeled by a multi-modal Gaussian, a weighted sum of the corresponding ground-truth interest distributions. To simulate a sequence of item interactions $\mathcal{V}_{u}$ (i.e., user history), we follow Mehta et al. (2023), first running a Markov Chain using a predefined user interest transition matrix to obtain the user’s interest interactions for $S=10$ steps. We do not consider the cold-start problem in this experiment, so we simply recommend one item from each generated cluster to form the user history (i.e., $|\mathcal{V}_{u}|=S$). The observation $\mathbf{o}_{u}$ over items in $\mathcal{V}_{u}$ is set to 1 if the item belongs to a ground-truth cluster, and -1 otherwise.
@@
User History and Observation Update. Using a predetermined user browsing model, clicked and skipped items are generated and appended to the user interaction history (and the corresponding user observation is likewise updated, 1 for clicked, -1 for skipped). This process continues until a maximum iteration of $T=10$ is reached. In this online setting, we adopt the dependent click model (DCM) of user browsing behavior, widely used in web search and recommendation (Cao et al., 2020; Chuklin et al., 2015). In the DCM, users begin by inspecting the top-ranked item, progressing down the list, engaging with items of interest and deciding to continue or terminate after each viewed item.

### v5

@@ 4.4 Item Embedding Pre-training
Following [2], we assume that item embeddings are fixed and precomputed: this ensures rapid computation and real-time updates at serving time. For item embedding pre-training, we use extreme multi-class classification [3, 29, 30]: given a training sample $(u_{i},v_{j})$, we first compute the likelihood of $u_{i}$ interacting with $v_{j}$, i.e., $p(v_{j}|u_{i})=\text{exp}(\mathbf{u}_{i}^{\intercal}\mathbf{v}_{j})/\sum_{v^{\prime}\in\mathcal{V}}\text{exp}(\mathbf{u}_{i}^{\intercal}\mathbf{v}^{\prime})$, where $\mathbf{u}_{i}$ and $\mathbf{v}_{j}$ are embeddings of $u_{i}$ and $v_{j}$. Our objective is to maximize the log-likelihood of a user interacting with their items. Moreover, we want to ensure that item embeddings align with their categories since categories explicitly indicate a user’s interests. We capture item-category information by computing the likelihood that an item belongs to a category in a similar way: $p(c_{k}|v_{j})=\text{exp}(\mathbf{v}_{j}^{\intercal}\mathbf{c}_{k})/\sum_{c^{\prime}\in\mathcal{C}}\text{exp}(\mathbf{v}_{j}^{\intercal}\mathbf{c}^{\prime})$. The overall objective for pre-training combines the two negative log-likelihoods using a scaling factor $\gamma$:
@@
where $\mathcal{V}_{u_{i}}$ is the set of items that $u_{i}$ interacted with and $\mathcal{C}_{v_{j}}$ is the set of categories that $v_{j}$ belongs to. The full item embedding matrix $\mathbf{V}$ is jointly learned and fixed after the pre-training phase. We do not use user and category embeddings after pre-training; other schemes, such as using item co-occurrence information to derive item embeddings without user or category embeddings, are possible. The full architecture for GPR4DUR is depicted in Figure 2.
4.5 Computational Complexity
The complexity of GPR is $O((\ell_{u})^{3})$, where $\ell_{u}$ is the number of training samples, due to covariance matrix inversion (Eq. 3 and Eq. 4). This inversion is required only once and is manageable with a few thousand examples. If the history length is too large, we either select a representative subset of interactions or focus on the most recent interactions (we adopt this latter approach in our experiments). Inference costs for each test point are $O(\ell_{u})$ for mean and $O((\ell_{u})^{2})$ for variance prediction, thus linear in the number of test points $|\mathcal{V}|$. As shown in Appendix A.7, our method has reasonable computational costs, slightly higher than strong baselines but with better performance on retrieval and ranking, demonstrating its applicability in real-world scenarios. Our approach is best-suited to scenarios with millions or tens of millions of items, when retrieving hundreds for ranking, as is common in real-world RSs (e.g., paper [31]). Similarly, in music recommendation, it would be especially useful for defining user interests over artists (a few millions at most per geographical region/language) vs. individual songs. Previous work improves GPR efficiency using Toeplitz and Kronecker structures, achieving $O(\ell_{u})$ training and $O(1)$ prediction complexity [32]. Other sublinear approximations partition item space to find candidates adaptively [33, 34]. While not our focus, it would be straightforward to apply these well-studied methods to our setting.
Finally we note that we do not expect users’ interest collections to be highly dynamic; hence, they will not require real-time updates. Our method is best-suited to use as a candidate generator, with the model updated periodically to create a pool of potentially interesting items for users.
@@ 5 Offline Experiments
We next evaluate our GPR4DUR on three real-world datasets, and compare our DUR technique with other state-of-the-art methods for multi-interest personalized retrieval.
@@
The higher the better $\Uparrow$ The higher the better $\Uparrow$ The lower the better $\Downarrow$ The higher the better $\Uparrow$
@@ 5.1 Datasets
We use three widely studied datasets: Amazon [35], MovieLens [16], and Taobao [36]. Following previous work [37, 38], we filter out items with fewer than 10 appearances and users with fewer than 25 interactions to ensure sufficient history for indicating multiple interests. Each interest corresponds to a category: for Amazon, we use the single principal category of each item; for MovieLens, we use 18 movie genres (each movie can belong to multiple genres); for Taobao, we cluster items into 20 categories using Ward clustering on the pretrained item embeddings as in [2]. The same categories are used during training and inference. Dataset statistics are shown in Table 6, and an ablation analysis of various clustering algorithms and numbers of clusters is provided in Table 8 (See Appendix).
@@ 5.2 Experiment Setup
Most prior work on recommendation and retrieval evaluates model performance on generalization to new items per user. Instead, similar to recent work [39, 3], we assess model performance on the ability to generalize to new users. We split users into disjoint subsets: training users ($\mathcal{U}^{\text{train}}$), validation users ($\mathcal{U}^{\text{val}}$), and test users ($\mathcal{U}^{\text{test}}$) in a ratio of 8:1:1. The last 20% of each user’s interaction sequence is treated as a holdout set for evaluation, and the first 80% as a history set for fitting the GPR model. We cap history length to 60, 160, and 100 for the three datasets, aligning with their average user interactions. For item embedding pre-training, we train the recommendation backbones using the history set of all training users and tune parameters based on the performance on the holdout set for validation users (details in Appendix A.3). Training is capped at 100,000 iterations with early stopping if validation performance does not improve for 50 successive iterations. We tune GPR hyperparameters by fitting the GP regressor to the history set of all training and validation users, then using the holdout set to fine-tune parameters (kernel and standard deviation). Metrics are reported on the holdout set for all test users (see Appendix A.5 for hyperparameter sensitivity).
@@ 5.3 Metrics
Metrics used in prior work on sequential recommendation are not well-suited to assess performance in our multi-interest retrieval task, for several reasons: (i) Conventional metrics, such as precision and recall, often employed in multi-interest research, do not adequately quantify whether an item list reflects the full range of a user’s (multiple) interests. A model may primarily recommend items from a narrow range of highly popular categories and still score high on these metrics while potentially overlooking niche interests. (ii) Metrics like precision and recall are overly stringent, only recognizing items in the recommendation list that appear in the holdout set. We argue that credit should also be given if a similar, though not identical, item is recommended (e.g., Iron Man 1 instead of Iron Man 2). This requires a soft version of these metrics. (iii) Multi-interest retrieval systems should expose users to niche content to address their diverse interests. However, conventional metrics may neglect the item perspective, potentially underserving users with specialized interests. Consequently, we propose the use of the following four metrics that encompassing the aforementioned factors.
@@
This metric is similar to subtopic-recall [40], which directly measures whether the model can comprehensively retrieve all user interests reflected in the holdout set. The higher the value of this metric the better:
@@
$\text{IR}@k$ $=\frac{1}{|\mathcal{U}^{\text{test}}|}\sum_{u\in\mathcal{U}^{\text{test}}}\frac{\sum\limits_{c\in\mathcal{C}(\mathcal{V}^{\text{d}}_{u})}\max\limits_{v_{i}\in\mathcal{V}^{\text{d}}_{u},v_{j}\in\mathcal{R}^{1:k}_{u}}S(v_{i},v_{j})}{|\mathcal{C}(\mathcal{V}^{\text{d}}_{u})|},\quad\text{s.t. }\mathcal{C}(v_{i})=\mathcal{C}(v_{j})=c,$ (7)
where $S(v_{i},v_{j})$ is the cosine similarity between item $v_{i}$ and $v_{j}$. To obtain ground-truth similarities between items, and to mitigate the influence of the chosen pre-trained model, we pretrain the item embeddings using YoutubeDNN [41] with a higher dimension size ($d=256$) to compute a uniform $\mathbf{S}_{i,j}$ for any backbone. A higher value of this metric is better.
@@
$\text{ED}@k$ $=\frac{1}{|\mathcal{U}^{\text{test}}|}\sum\limits_{u\in\mathcal{U}^{\text{test}}}||\epsilon^{*}_{u}-\epsilon_{u}^{1:k}||_{2}^{2},\quad\text{s.t., }\epsilon^{*}_{u,c}=\frac{\sum_{v\in\mathcal{V}^{\text{d}}_{u}}\mathbbm{1}_{{c\in\mathcal{C}(v)}}}{\sum_{v\in\mathcal{V}^{\text{d}}_{u}}|\mathcal{C}(v)|},\epsilon^{1:k}_{u,c}=\frac{\sum_{v\in\mathcal{R}^{1:k}_{u}}\mathbbm{1}_{{c\in\mathcal{C}(v)}}}{\sum_{v\in\mathcal{R}^{1:k}_{u}}|\mathcal{C}(v)|}.$ (8)
@@ 5.4 Methods Studied
Our overall performance comparison addresses two central research questions (RQs): whether our proposed method offers superior retrieval performance for users with multiple interests (RQ1); and whether it induces appropriate item-sided exposure w.r.t. both popular and niche interests (RQ2).
Performance on the retrieval task. To assess RQ1, we evaluate the effectiveness of the retrieval phase. As shown in Table 1, GPR4DUR outperforms almost all baselines across datasets w.r.t. Interest Coverage (IC@$k$) and Interest Relevance (IR@$k$), demonstrating high degrees of retrieval coverage and relevance. Specifically, on the Amazon dataset, GPR4DUR achieves the highest performance in 4 out of 6 interest metrics, in MovieLens it leads in 5 out of 6 metrics, and in Taobao, it excels across all 6 interest metrics. This shows that GPR4DUR covers a wide range of user interests while simultaneously maintaining high relevance. To assess the statistical significance of our model’s performance compared to the best baseline, we performed a paired t-test, which evaluates whether the means of two paired samples differ significantly, ensuring that the observed improvements are not due to random variation.
For RQ2, our objective is to ascertain whether item exposure is suitably balanced. We measure this using the Exposure Deviation (ED@$k$) and Tail Exposure Improvement (TEI@$k$) metrics. Lower values of ED@$k$ suggest more satisfying category exposure, while higher values of TEI@$k$ are indicative of enhanced exposure in the long tail item categories. GPR4DUR is highly effective on these metrics, consistently performing well across all datasets in most cases, and validating its ability to provide an optimal level of category exposure. Notably, GPR4DUR achieves the best results in 8 of 9 TEI@$k$ metrics, indicating its superior ability to generate exposure to niche categories/interests. We point out, however, that all TEI@$k$ values are negative, which suggests that none of the methods we tested improve exposure for niche categories relative to the exposure in the user holdout sets. This is explained by the well-known popularity bias inherent in most recommendation methods, and suggests that further work is required on diversification strategies to mitigate such effects.
Performance on the ranking task. While our primary emphasis is on the retrieval phase, we also present model comparisons for the ranking task. As illustrated in Table 2, GPR4DUR demonstrates competitive performance on the traditional relevance metrics relative to the baseline models. This further validates the strength of our proposed method.
@@
(RQ4) We underscored the importance of efficiency for a good user representation in Sec. 1. To shed light on this, we examine the IC@20 and IR@20 performance of various methods on Amazon. As shown in Fig. 4, GPR4DUR is effective w.r.t. interest coverage even when operating with low-dimensional embeddings (i.e., $d=8$ and $d=4$). By contrast, the performance of the SUR and MUR methods degrades as the dimensionality decreases. GPR4DUR consistently outperforms other methods w.r.t. interest relevance across all dimensionalities examined. We note that lower dimensions facilitate higher interest relevance due to increased cosine similarity (Eq. 7), explaining the inverse correlation between IR@20 and dimension size in Figure 4 (right).
@@ 6 Online Simulation
Data Preparation. We assume $|\mathcal{C}|=10$ interest clusters, each represented by a $d$-dim ($d=32$) multivariate Gaussian. We randomly select a (user-specific) subset of these interests as the ground-truth interest set for each user. We set $|\mathcal{U}|=1000$ and $|\mathcal{V}|=3000$, with $300$ items in each interest cluster. Each item belongs to a single cluster and its embedding is sampled from the corresponding interest distribution. Each user is modeled by a multi-modal Gaussian, a weighted sum of the corresponding ground-truth interest distributions. To simulate a sequence of item interactions $\mathcal{V}_{u}$ (i.e., user history), we follow [46], first running a Markov Chain using a predefined user interest transition matrix to obtain the user’s interest interactions for $S=10$ steps. We do not consider the cold-start problem in this experiment, so we simply recommend one item from each generated cluster to form the user history (i.e., $|\mathcal{V}_{u}|=S$). The observation $\mathbf{o}_{u}$ over items in $\mathcal{V}_{u}$ is set to 1 if the item belongs to a ground-truth cluster, and -1 otherwise.
@@
User History and Observation Update. Using a predetermined user browsing model, clicked and skipped items are generated and appended to the user interaction history (and the corresponding user observation is likewise updated, 1 for clicked, -1 for skipped). This process continues until a maximum iteration of $T=10$ is reached. In this online setting, we adopt the dependent click model (DCM) of user browsing behavior, widely used in web search and recommendation [47, 48]. In the DCM, users begin by inspecting the top-ranked item, progressing down the list, engaging with items of interest and deciding to continue or terminate after each viewed item.